OpenStudy (lovedance11):

Subtract. 7 - 5 - - 16 16

Mehek (mehek14):

\(\dfrac{7}{16}-\dfrac{5}{16}\) ?

OpenStudy (lovedance11):

@Mehek14 What is the answer?

Mehek (mehek14):

just subtract 7 - 5

OpenStudy (lovedance11):

ok

Mehek (mehek14):

\(\dfrac{7-5}{16}\)